OpenStudy (anonymous):

Use linear approximation to estimate 15^.25

OpenStudy (anonymous):

Or, in other words, use linear estimation to estimate \[\sqrt[4]{15}\]

OpenStudy (anonymous):

well \[16^{1/4}=2\] y(x)=x^1/4 y'(x)= 1/4 x^-3/4 \[y(x+\Delta)=y(x)+y'(x)\Delta x\]

OpenStudy (anonymous):

To be honest, I'm kind of confused as to what exactly linear approximation is. I know that the fourth root of 15 is an irrational number; is it essentially a pen and paper method for approximating its value to a relevant number of significant figures?

OpenStudy (anonymous):

And thank you for your help so far, by the way.

OpenStudy (anonymous):

2+(1/32)(-1)=1.97

OpenStudy (anonymous):

Actually it returns pretty accurate number provided that you approximate closed to known value(in our case 16)

OpenStudy (anonymous):

Ah, I understand your mathematical logic thus far. So \[\Delta\]x in this case is -1 because you're taking the rate of change at that point in the function and accounting for the change in x-value relative to the y value you originally took (16)?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Cool, thanks guys.

OpenStudy (dean.shyy):

This may be helpful to you: http://is.gd/8MFOpZ